Find an equation of a parabola with
a) focus(0,-7) and directrix x = 7
b) focus(5,0) and directrix x = -5
Thanks so much in advance:)

Answer by lwsshak3(11628)

Find an equation of a parabola with
a) focus(0,-7) and directrix: x=7
This is an equation of a parabola that opens left with vertex at origin.
Its basic form of equation: y^2=-4px
For given parabola:
vertex:(0,0)
axis of symmetry: y=0
p=7 (distance from vertex to directrix and focus on the axis of symmetry)
4p=28
equation: y^2=-28x
...
b) focus(5,0) and directrix x = -5
This is an equation of a parabola that opens right with vertex at origin.
Its basic form of equation: y^2=4px
For given parabola:
vertex:(0,0)
axis of symmetry: y=0
p=5 (distance from vertex to directrix and focus on the axis of symmetry)
4p=20
equation: y^2=20x